Either of the cost alternatives shown below can be used in a chemical refining process. If the company’s MARR is 15% per year, determine which should be selected on the basis of an incremental ROR analysis.

A

B

First cost ,$

40,000

61,000

Annual cost, $/year

25,000

19,000

Salvage value, $

8,000

11,000

Life, years 5

5

5


      -       A.       B.       C.       D.       E.       F.   

The incremental rate of return computed using a present worth analysis is nearest to is the incremental rate of return to be found

      -       A.       B.       C.       D.       E.       F.   

Which alternative is to be chosen?

A.

15%

B.

17%

C.

Alternative A

D.

16%

E.

14%

F.

Alternative B
Answer #1

MARR = 15%

Incremental cost (B-A) = 61000 - 40000 = 21000

Annual saving = -19000 + 25000 = 6000

Incremental salvage value = 11000 - 8000 = 3000

t = 5 yrs

Let incremental rate of return be i%, then PW at i% =0

PW = -21000 + 6000*(P/A,i%,5) + 3000*(P/F,i%,5) = 0

6000*(P/A,i%,5) + 3000*(P/F,i%,5) = 21000

dividing by 3000

2*(P/A,i%,5) + (P/F,i%,5) = 7

Using trail and error method

When i = 15%, 2*(P/A,i%,5) + (P/F,i%,5) = 7.20149

When i = 16%, 2*(P/A,i%,5) + (P/F,i%,5) = 7.02470

When i = 17%, 2*(P/A,i%,5) + (P/F,i%,5) = 6.85480

using interpolation

i = 16% + (7.02470-7) / (7.02470-6.85480) *(17-16)

i = 16% + 0.14537

i = 16.15% (Approx) (option D)

As Incremental IRR is greater than MARR, Project B should be selected (option F)